A particle is moving with the given data. Find the position of the particle.

v(t) = 1.5√t
s(4) = 12

s(t) =
1
Expert's answer
2011-03-17T06:57:04-0400
s(t ) = int(v(t), dt) = 3/2 * 2/3 t3/2 + 4 = t3/2 + 4